OpenStudy (anonymous):

integrate of sin5x/sec^2x dx

OpenStudy (kc_kennylau):

\[\int\frac{\sin5x}{\sec^2x}dx\]

OpenStudy (kc_kennylau):

Is this what you mean?

OpenStudy (anonymous):

yeah

OpenStudy (kc_kennylau):

@ganeshie8 @zepdrix

zepdrix (zepdrix):

You can get the sin5x in terms of sinx and cosx by applying the Angle Addition Formula a billion times.. there must be a better way though... hmm

OpenStudy (kc_kennylau):

I'm also looking for a better way lol

OpenStudy (aravindg):

"billion times" :O

ganeshie8 (ganeshie8):

\(\large \int\frac{\sin5x}{\sec^2x}dx \) \(\large \frac{1}{2}\int \sin 5x (1+\cos 2x)dx \) \(\large \frac{1}{2}\int \sin 5x dx + \frac{1}{2}\int \sin (5x)\cos (2x)dx \) \(\large \frac{1}{2}\int \sin 5x dx + \frac{1}{4}\int (\sin(7x) + \sin(3x))dx \)

OpenStudy (anonymous):

\[\int\limits \frac{ \sin ^5x }{\sec ^2x }dx=\int\limits \sin ^5x \left( 1-\sin ^2x \right)dx\] \[\int\limits \sin ^5x~dx-\int\limits \sin ^7x~dx\] \[\int\limits \sin ^5x~dx=\int\limits \sin ^4x~\sin x~dx\] \[=\int\limits \left( 1-\cos ^2x \right)^2\sin x~dx=\int\limits \left( 1-2\cos ^2x+\cos ^4x \right)\sin x~dx\] you can complete now. \[\int\limits f^n(x)f \prime \left( x \right)dx=\frac{ f ^{n+1}(x) }{n+1}\]

OpenStudy (anonymous):

if it is sin 5x then ganeshie8 solution is best